OpenStudy (annipuppi):

Help me find the circumcenter of the triangle with the coordinates: B(2,2) P(4,10) H(12,2)

OpenStudy (annipuppi):

@ganeshie8

OpenStudy (anonymous):

(G+H)/2 = (1+6, 0+0)/2 = (7/2, 0) Because it's along the x-axis, (slope, m=0) its perpendiculars all have equations, x = c, which are all parallel to the y-axis (and with slope = -1/m = ∞). Here, it's x = 7/2 Taking GJ next, its midpoint is (G+J)/2 = (1+3, 0-6)/2 = (2, -3) Slope of GJ is given by m = (-6-0)/(3-1) = -6/2 = -3 Slope of the perp is a = -1/m = 1/3 So we're looking for a line of slope = 1/3 passing through (2, -3): y = x/3 + b --> -3 = 2/3 + b b = -11/3 y = x/3 - 11/3 Together with our first equation, x = 7/2, gives y = 7/6 - 11/3 = -15/6 = -5/2 So (x,y) = (7/2, -5/2) The third P.B., of HJ, its midpoint is (H+J)/2 = (9/2, -3) Slope of HJ = (0-(-6))/(6-3) = 6/3 = 2 Slope of P.B. of HJ = -1/2; P.B.passes through (9/2, -3): y = -x/2 + b -3 = -9/4 + b; b = -3/4 y = -x/2 - 3/4 To check that the point we already found lies on this line, we just plug the point's coordinates in: -5/2 =? -(7/2)/2 - 3/4 = -7/4 - 3/4 = /10/4 = -5/2 Success!

OpenStudy (annipuppi):

My triangle doesn't have a G.. I'm confused @zobobozo

ganeshie8 (ganeshie8):

you're here @annipuppi ?

OpenStudy (annipuppi):

Now I am @ganeshie8

OpenStudy (rational):

hey

OpenStudy (rational):

draw it out

OpenStudy (rational):



OpenStudy (annipuppi):

Yeah I've done that but I don't get it still.